### v4

@@ 7 Discussion
This paper presents , a new distortion aware vision transformer which adapts its structure to the lens distortion profile of a (calibrated) lens. achieves state-of-the-art performance on zero-shot adaptation (without pretraining) on different lenses on classification using synthetically distorted images from the ImageNet1k dataset.
While the original idea was introduced in [26], this paper presents three new contributions, significantly extending the encoder-only architecture. First, a self-calibration network is used to calibrate the lens profile, resulting in an entirely uncalibrated pipeline for wide-angle image recognition. Second, DarSwin-Unet is an extension of encoder to a full encoder-decoder distortion-aware architecture. Third, we demonstrate that DarSwin-Unet achieves state-of-the-art performance on zero-shot adaptation (without pretraining) on different lenses for per-pixel depth estimation using synthetically distorted images from the Matterport3D datasets, respectively.
@@
Despite strong generalization capabilities, our method still has limitations that should be explored in future work. First, while the distortion-aware sampling strategy was shown to be effective for zero-shot adaptation, the sparsity of sampling points and the necessity to interpolate pixel values may affect the performance of the model. While this issue is partially alleviated using our proposed jittering augmentation technique, other strategies may also be possible. This is also a problem in per-pixel tasks such as depth estimation, which must rely on a $k$-NN for mapping the samples back to pixel space. Second, our model assumes knowledge of the lens distortion profile; hence, it is appropriate only for the calibrated case. For the classification experiment, as shown in the above section, we took inspiration from [72] to calibrate the uncalibrated lens using a network. However, the limitation persists for the depth estimation task.

### v5

@@ 6 Discussion
This paper presents DarSwin, a new distortion aware vision transformer which adapts its structure to the lens distortion profile of a (calibrated) lens. DarSwin achieves state-of-the-art performance on zero-shot adaptation (without pretraining) on different lenses on classification using synthetically distorted images from the ImageNet1k dataset.
@@
While our method demostrates state-of-the-art performance, it suffers from some limitations. First, the distortion aware sampling strategy is shown to be effective for zero-shot adaptation, the sparsity of sampling points and necessity to interpolate pixel values may affect the performance of the model. While this issue is partially alleviated using our proposed jittering augmentation technique, other strategies may also be possible. Second, our model assumes knowledge of the lens distortion profile, hence it is appropriate only for the calibrated case. We hope to extend our work to uncalibrated lenses, for example by taking inspiration from [15, 47, 10]. Finally, while our experiments demonstrate promising performance on classification experiments, we wish to expand to per-pixel tasks, such as semantic classification and depth estimation. Here, making pixel decoders distortion aware is an exciting direction for future work.
Acknowledgments
This research was supported by NSERC grant ALLRP-567654, Thales, an NSERC USRA to J. Lagüe, and the Digital Research Alliance Canada. We thank Yohan Poirier-Ginter, Frédéric Fortier-Chouinard, Adam Tupper and Justine Giroux for proofreading.